# VISTA: Enhancing Vision-Text Alignment in MLLMs via Cross-Modal Mutual Information Maximization

### v1

@@ C.2 General multimodal benchmarks
MMBench MMBench(Multimodal Benchmark)[32] is a task-driven benchmark that focuses on systematically evaluating multimodal models across diverse real-world application scenarios, such as visual question answering, image captioning, and video understanding. Its emphasis on practical use cases highlights its importance for assessing the practical utility of MLLMs.
SEED: SEED(Spatial and Entity-aware Evaluation Dataset)[33] is a benchmark specifically designed to evaluate the spatial and entity reasoning capabilities of multimodal models. By incorporating complex spatial relationships and entity-based queries, SEED tests a model’s ability to perform fine-grained reasoning, which is critical for tasks such as scene understanding and object-oriented question answering.
AI2D: AI2D(Allen Institute for AI Diagram Dataset)[34] is a dataset centered on diagram understanding, designed to evaluate models’ abilities to process non-photographic visual content. It focuses on reasoning over diagrams and charts, making it vital for tasks requiring scientific and technical visual comprehension.
MMMU: MMMU(Multimodal Multitasking Understanding)[35] evaluates the multitasking capabilities of multimodal models by testing their performance on multiple simultaneous tasks across different modalities. This benchmark is essential for assessing the adaptability and efficiency of models in dynamic, multitask scenarios.
MMTB: MMTB(Multimodal Task Benchmark)[36] is a broad benchmark designed to evaluate the performance of multimodal models on a wide range of tasks, including vision-and-language navigation, multimodal reasoning, and image captioning. Its diversity makes it a strong indicator of a model’s overall multimodal proficiency.
OCRB: OCRB (Optical Character Recognition Benchmark)[24] is a specialized benchmark for assessing a model’s ability to recognize and interpret text in images. It focuses on OCR-related tasks, such as text detection, transcription, and contextual understanding, which are crucial for applications like document analysis and scene-text understanding.
MMStar: MMStar[37], an elite vision-indispensable multi-modal benchmark comprising 1,500 samples meticulously selected by humans. MMStar benchmarks 6 core capabilities and 18 detailed axes, aiming to evaluate LVLMs’ multi-modal capacities with carefully balanced and purified samples.
@@ C.3 Fine-grained Visual Perception and Referring Expression Comprehension Benchmarks
MME: [38] (Multimodal Evaluation) is a comprehensive benchmark designed to assess a model’s proficiency in understanding and reasoning across multiple modalities, including vision, language, and audio. Comprising 14 diverse subtasks, MME evaluates both perceptual and cognitive aspects of multimodal intelligence, providing a standardized framework to measure cross-modal reasoning and generalization capabilities of multimodal large language models (MLLMs).
RefCOCO: [39] is a widely adopted referring expression comprehension dataset that challenges models to localize objects within images based on natural language descriptions. It emphasizes the alignment between fine-grained visual features and linguistic cues in everyday scenes, serving as a fundamental testbed for grounding language in vision.
RefCOCO+: [39] builds upon RefCOCO by excluding absolute spatial references (e.g., "left", "right") from the expressions, thereby compelling models to rely predominantly on appearance-based attributes for object localization. This constraint elevates the difficulty of the task and probes the model’s ability to discern subtle visual details.
RefCOCO+: [39] builds upon RefCOCO by excluding absolute spatial references (e.g., "left", "right") from the expressions, thereby compelling models to rely predominantly on appearance-based attributes for object localization. This constraint elevates the difficulty of the task and probes the model’s ability to discern subtle visual details.
RefCOCOg: [40] extends the complexity of referring expressions by providing longer, more descriptive, and contextually rich language annotations. This dataset evaluates a model’s capacity to integrate detailed linguistic information with visual content, reflecting more naturalistic and challenging scenarios for referring expression comprehension.
@@ C.4 Lmms Eval
We utilize lmms eval (v0.3.0) [42] as the primary evaluation framework for our benchmark. lmms eval is a highly extensible and modular platform designed to accommodate a wide spectrum of evaluation tasks for multimodal large models. At its core, the framework leverages a declarative YAML configuration format, enabling precise specification and reproducibility of evaluation protocols. This design not only facilitates straightforward task definition—such as swapping dataset paths to evaluate on new benchmarks—but also supports sophisticated customization of evaluation logic, empowering researchers to tailor assessments to nuanced experimental setups. By coupling configuration files with explicit codebase commit hashes, lmms eval ensures that evaluation environments can be reliably shared and replicated across the community, thereby promoting transparency and rigor in model assessment. This combination of flexibility, reproducibility, and ease of use makes lmms eval an indispensable tool for advancing standardized evaluation in multimodal research.
@@ Appendix D Experiments Detail
We adopt hyperparameter configurations consistent with those reported in TinyLLaVA and LLaVA to ensure comparability. Notably, our TinyLLaVA [43] baseline results are derived from our own experimental runs rather than relying on official benchmarks. Although TinyLLaVA  originally utilizes the ShareGPT4V dataset, we exclusively employ the LLaVA[27] dataset in all experiments. This deliberate choice is driven by the need to maintain dataset uniformity, thereby enabling a controlled investigation of model scaling effects on performance without confounding variables introduced by differing data distributions.

### v2

@@ C.2 General multimodal benchmarks
MMBench MMBench(Multimodal Benchmark)Liu et al. [2025] is a task-driven benchmark that focuses on systematically evaluating multimodal models across diverse real-world application scenarios, such as visual question answering, image captioning, and video understanding. Its emphasis on practical use cases highlights its importance for assessing the practical utility of MLLMs.
SEED: SEED(Spatial and Entity-aware Evaluation Dataset)Li et al. [2023a] is a benchmark specifically designed to evaluate the spatial and entity reasoning capabilities of multimodal models. By incorporating complex spatial relationships and entity-based queries, SEED tests a model’s ability to perform fine-grained reasoning, which is critical for tasks such as scene understanding and object-oriented question answering.
AI2D: AI2D(Allen Institute for AI Diagram Dataset)Kembhavi et al. [2016] is a dataset centered on diagram understanding, designed to evaluate models’ abilities to process non-photographic visual content. It focuses on reasoning over diagrams and charts, making it vital for tasks requiring scientific and technical visual comprehension.
MMMU: MMMU(Multimodal Multitasking Understanding)Yue et al. [2024] evaluates the multitasking capabilities of multimodal models by testing their performance on multiple simultaneous tasks across different modalities. This benchmark is essential for assessing the adaptability and efficiency of models in dynamic, multitask scenarios.
MMTB: MMTB(Multimodal Task Benchmark)Ying et al. [2024] is a broad benchmark designed to evaluate the performance of multimodal models on a wide range of tasks, including vision-and-language navigation, multimodal reasoning, and image captioning. Its diversity makes it a strong indicator of a model’s overall multimodal proficiency.
OCRB: OCRB (Optical Character Recognition Benchmark)Mishra et al. [2019] is a specialized benchmark for assessing a model’s ability to recognize and interpret text in images. It focuses on OCR-related tasks, such as text detection, transcription, and contextual understanding, which are crucial for applications like document analysis and scene-text understanding.
MMStar: MMStarChen et al. [2024a], an elite vision-indispensable multi-modal benchmark comprising 1,500 samples meticulously selected by humans. MMStar benchmarks 6 core capabilities and 18 detailed axes, aiming to evaluate LVLMs’ multi-modal capacities with carefully balanced and purified samples.
@@ C.3 Fine-grained Visual Perception and Referring Expression Comprehension Benchmarks
MME: Fu et al. [2024] (Multimodal Evaluation) is a comprehensive benchmark designed to assess a model’s proficiency in understanding and reasoning across multiple modalities, including vision, language, and audio. Comprising 14 diverse subtasks, MME evaluates both perceptual and cognitive aspects of multimodal intelligence, providing a standardized framework to measure cross-modal reasoning and generalization capabilities of multimodal large language models (MLLMs).
RefCOCO: Kazemzadeh et al. [2014] is a widely adopted referring expression comprehension dataset that challenges models to localize objects within images based on natural language descriptions. It emphasizes the alignment between fine-grained visual features and linguistic cues in everyday scenes, serving as a fundamental testbed for grounding language in vision.
RefCOCO+: Kazemzadeh et al. [2014] builds upon RefCOCO by excluding absolute spatial references (e.g., ”left”, ”right”) from the expressions, thereby compelling models to rely predominantly on appearance-based attributes for object localization. This constraint elevates the difficulty of the task and probes the model’s ability to discern subtle visual details.
RefCOCO+: Kazemzadeh et al. [2014] builds upon RefCOCO by excluding absolute spatial references (e.g., ”left”, ”right”) from the expressions, thereby compelling models to rely predominantly on appearance-based attributes for object localization. This constraint elevates the difficulty of the task and probes the model’s ability to discern subtle visual details.
RefCOCOg: Mao et al. [2016] extends the complexity of referring expressions by providing longer, more descriptive, and contextually rich language annotations. This dataset evaluates a model’s capacity to integrate detailed linguistic information with visual content, reflecting more naturalistic and challenging scenarios for referring expression comprehension.
@@ C.4 Lmms Eval
We utilize lmms eval (v0.3.0) Li et al. [2024] as the primary evaluation framework for our benchmark. lmms eval is a highly extensible and modular platform designed to accommodate a wide spectrum of evaluation tasks for multimodal large models. At its core, the framework leverages a declarative YAML configuration format, enabling precise specification and reproducibility of evaluation protocols. This design not only facilitates straightforward task definition—such as swapping dataset paths to evaluate on new benchmarks—but also supports sophisticated customization of evaluation logic, empowering researchers to tailor assessments to nuanced experimental setups. By coupling configuration files with explicit codebase commit hashes, lmms eval ensures that evaluation environments can be reliably shared and replicated across the community, thereby promoting transparency and rigor in model assessment. This combination of flexibility, reproducibility, and ease of use makes lmms eval an indispensable tool for advancing standardized evaluation in multimodal research.
@@ Appendix D Experiments Detail
We adopt hyperparameter configurations consistent with those reported in TinyLLaVA and LLaVA to ensure comparability. Notably, our TinyLLaVA Zhou et al. [2024] baseline results are derived from our own experimental runs rather than relying on official benchmarks. Although TinyLLaVA  originally utilizes the ShareGPT4V dataset, we exclusively employ the LLaVALiu et al. [2024a] dataset in all experiments. This deliberate choice is driven by the need to maintain dataset uniformity, thereby enabling a controlled investigation of model scaling effects on performance without confounding variables introduced by differing data distributions.